OpenStudy (anonymous):

Use the word SCHOOL to answer this question. If the letters of this word are written on a paper and then cut into squares with one letter per square, what is the probability of the following event: P(H or L)?

OpenStudy (across):

In the word \(\text{SCHOOL}\), there is a total of \(6\) letters. Moreover, the letters \(\text{H}\) and \(\text{L}\) appear only once. Therefore, the probability that you randomly pick one of them is \(1/6\). As a result, \(P(\text{H or L})=P(\text{H})+P(\text{L})-P(\text{H and L})\).

OpenStudy (anonymous):

Alright, Thank you.

OpenStudy (anonymous):

It was incorrect.

hartnn (hartnn):

what was incorrect, what value did u put ?

OpenStudy (anonymous):

I clicked 1/6 and it said it was incorrect.

hartnn (hartnn):

because its not 1/6! as @across explained u P(H)=P(L)=1/6

OpenStudy (anonymous):

Yeah :/

hartnn (hartnn):

but u need P(H or L) = 1/6+1/6-0 = ??

OpenStudy (across):

Actually, here's another hint: \(P(\text{H and L})=P(\text{H})P(\text{L})\).

OpenStudy (anonymous):

What do you mean hartnn?

hartnn (hartnn):

but H and L can't be picked together, so P(H and L) = 0 isn't it ?

OpenStudy (across):

You are right; I forgot we are only picking one letter, not two in sequence. :)

OpenStudy (anonymous):

My options are; 1/36, 1/3,1/6, and 1 but 1/6 is incorrect.

OpenStudy (across):

@RebelChi, @hartnn broke it down for you:\[\frac16+\frac16=?\]

OpenStudy (anonymous):

1/36?

hartnn (hartnn):

is there + sign or * sign in between @RebelChi ? how 1/36 ?

OpenStudy (anonymous):

What do you mean?

hartnn (hartnn):

this is basics: \(\huge \frac{1}{6}+\frac{1}{6}=\frac{1+1}{6}=\frac{2}{6}=\frac{1}{3}\)

OpenStudy (anonymous):

Ohh so you have to simplify it?

hartnn (hartnn):

yes, and on simplification it gives 1/3.

OpenStudy (anonymous):

Okay, I'll click it and see if it's right. :)

OpenStudy (anonymous):

Correct! Yay :P

hartnn (hartnn):

great :)